# supplement_arXiv

### v2

@@ Abstract
This paper introduces a robust approach to functional principal component analysis (FPCA) for compositional data, particularly density functions. While recent papers have studied density data within the Bayes space framework, there has been limited focus on developing robust methods to effectively handle anomalous observations and large noise. To address this, we extend the Mahalanobis distance concept to Bayes spaces, proposing its regularized version that accounts for the constraints inherent in density data. Based on this extension, we introduce a new method, robust density principal component analysis (RDPCA), for more accurate estimation of functional principal components in the presence of outliers. The method’s performance is validated through simulations and real-world applications, showing its ability to improve covariance estimation and principal component analysis compared to traditional methods.
@@ 1 Introduction
In the recent decade, functional data analysis (FDA) has grown rapidly (e.g., Ramsay and Silverman (2005)), but there has been less focus on functional distributional data, like density functions. In this context, we do not just understand probability density functions by densities but rather consider a broader class containing relative data that has a functional nature. Examples are age-specific fertility and mortality rates or spectra analysis of different materials. In all these cases, the shape of the underlying curve is of interest, rather than the absolute values taken by the functions.
Densities as random objects need special handling because they are non-negative functions that carry relative information and can be presented with unit integral constraints. This makes compositional data analysis (CoDA) a good framework for densities. By extending CoDA principles to infinite dimensions, densities can be thought of as “infinite-dimensional compositions” and analyzed within a particular Hilbert space called Bayes spaces (Egozcue et al., 2006; van den Boogaart et al., 2010).
Functional principal component analysis (FPCA), adapted for Bayes spaces as simplicial functional PCA (SFPCA) (Hron et al., 2016), helps to analyze density data, but its reliance on the sample covariance makes it sensitive to outliers. To address this, a trimmed covariance estimator can be used instead, where center-outwards ordering is used to focus on typical observations. Depth measures provide one possible way to order data that could improve robustness (Fraiman and Muniz (2001); Febrero et al. (2008); López-Pintado and Romo (2009); Arribas-Gil and Romo (2014); Chakraborty and Chaudhuri (2014)). For instance, “spatial depth” has already been useful for data in Bayes spaces Menafoglio et al. (2021). Alternatively, the Mahalanobis distance has been traditionally used for this purpose, and it has been extended to high-dimensional multivariate settings (Secchi et al., 2013), functional settings by, e.g., Galeano et al. (2015); Ghiglietti et al. (2017); Berrendero et al. (2020) for robust covariance estimation, as in the minimum regularized covariance trace (MRCT) estimator (Oguamalam et al., 2024).
The aim of this study is two-fold: first, we extend the concept of regularized Mahalanobis distance for univariate functional data to Bayes spaces, which we will refer to as RDMD. Using the defined notion of Mahalanobis distance, we obtain a robust covariance estimator that further leads to a robust PCA method specifically designed for density data. The resulting procedure is able to handle the density constraints and reduce outlier sensitivity for infinite dimensional compositions.
@@
The rest of the paper is organized as follows: Section 2 discusses the concept of Bayes spaces and SFPCA and showcases potential problems that the outliers in the sample can cause. For an adaptation into the Bayes Hilbert spaces, concepts of data standardization and consequently, the notion of Mahalanobis distance are crucial. In Section 3 we, therefore, propose a way of standardizing (whitening) relative functional data, thus extending and generalizing the notions of truncated- and $\alpha$-Mahalanobis distance from Galeano et al. (2015) and Berrendero et al. (2020), respectively to Bayes Hilbert spaces. In Section 4, the novel, robust density PCA (RDPCA) method is introduced. Its performance is evaluated in a simulation study and real-data examples in Sections 5 and 6, respectively. Sections LABEL:supp:simulations and LABEL:supp:realdata in the supplement provide some additional results, respectively. The final Section 7 concludes with a discussion. Proofs of technical results are given in Supplement Section LABEL:supp:proofs.
@@ 2.1 Bayes space
If we expand the horizon to infinite dimensional compositions, we consider positive functions that are typically represented with unit integral constraint. Intuitively, such compositions can be thought of as density functions. Therefore, statistical concepts designed for unconstrained functional data need to be adapted, as they do not account for this feature, leading to potentially meaningless results when applied to densities (Delicado, 2011; Hron et al., 2016; Menafoglio et al., 2018). Consequently, to work with functional compositions, it is necessary to establish the fundamental concepts of the underlying space structure.
The foundational results were initially presented by Egozcue et al. (2006) and were subsequently extended by van den Boogaart et al. (2010), van den Boogaart et al. (2014) and Talská et al. (2020). The framework for probability density functions is based on a measurable space $(\Omega,\mathcal{A})$, typically set to $(\mathbb{R},\mathfrak{B}(\mathbb{R}))$, with $\mathfrak{B}(\mathbb{R})$ as the Borel $\sigma$-algebra on $\mathbb{R}$. Within this space, we consider the set of measures $\mu$ that are absolutely continuous with respect to a reference measure $\lambda$, here assumed to be the Lebesgue measure with compact support $I=[a,b]\subset\mathbb{R}$. While the authors of van den Boogaart et al. (2014) developed the theory of Bayes space for more general measures $\lambda$, in practice, commonly the Lebesgue measure is chosen. Furthermore, usually the measures $\mu$ are identified with their respective Radon-Nikodym derivative, density $f$. For these densities, a notion of equivalence in the Bayes sense for multiples is introduced, i.e $f=_{B}g$ if $f=c\cdot g$ for some $c\in\mathbb{R}.$ W.l.o.g. within these $(=_{B})$-equivalence classes, we take the element satisfying the unit integral constraint as the representative.
[…]
$(f\oplus g)(t)=_{B}f(t)g(t),\quad(\alpha\odot f)(t)=_{B}(f(t))^{\alpha},\quad\langle f,g\rangle_{B^{2}}=_{B}\int_{I}\int_{I}\text{log}\frac{f(t)}{f(s)}\text{log}\frac{g(t)}{g(s)}\mathrm{d}t\,\mathrm{d}s,$
for densities $f,g$ and $\alpha\in\mathbb{R}$. Endowing the subspace $\mathcal{B}^{2}(I)$ of positive densities on $I$ with a square-integrable logarithm (van den Boogaart et al., 2014) with this structure yields a separable Hilbert space. In practice, the centered logratio (clr) transformation
$\text{clr}(f)=\text{log}(f)-\frac{1}{b-a}\int_{I}(\text{log}(f))(t)\mathrm{d}t$ (1)
@@
$\langle f,g\rangle_{B^{2}}=~{}\langle\text{clr}(f),\text{clr}(g)\rangle_{L^{2}}~{}~{}\text{and}~{}~{}\text{clr}((\alpha\odot f)\oplus(\beta\odot g))=~{}\alpha\cdot\text{clr}(f)+\beta\cdot\text{clr}(g).$
@@ 2.2 Moments of random densities
for every $f,g\in\mathcal{B}^{2}(I)$. We write
$C_{B^{2},X}=\mathbb{E}((X\ominus\mu_{B^{2},X})\otimes_{B^{2}}(X\ominus\mu_{B^{2},X})),$
where for $f,g\in\mathcal{B}^{2}(I)$ the outer product $f\otimes_{B^{2}}g:\mathcal{B}^{2}(I)\to\mathcal{B}^{2}(I)$ is the linear operator satisfying $(f\otimes_{B^{2}}g)(h)=\langle g,h\rangle_{B^{2}}f$, for $h\in\mathcal{B}^{2}(I)$.
Lemma 2.1 gives the connection between the Bayes mean density $\mu_{B^{2},X}$ and covariance $C_{B^{2},X}$ and the “traditional" $L^{2}$-mean and the covariance of $\mathrm{clr}(X)$. The proof of Lemma 2.1 can be found in Section (LABEL:supp:proofs) of the supplement.
@@ 2.3 Simplicial functional PCA
As in the multivariate principal component analysis, functional PCA aims to search for the main modes of variability within a dataset, see e.g. Ramsay and Silverman (2005). More recently, this concept was extended to the Bayes spaces setting in Hron et al. (2016). The authors coined the term simplicial functional PCA referring to the simplex, which is the multivariate counterpart of the Bayes space. More precisely, given a data set $X_{1},\ldots,X_{n}\in\mathcal{B}^{2}(I)$, the problem of finding the first principal function $\zeta_{1}$ is defined as
$\underset{\zeta\in\mathcal{B}^{2}(I)}{\text{max}}\frac{1}{n}\sum_{i=1}^{n}\langle X_{i}\ominus\hat{\mu}_{B^{2},X},\zeta\rangle_{B}^{2}~{}~{}~{}\text{subject to~{}}||\zeta||_{B}^{2}=1,$ (3)
@@
However, as previously emphasized, the eigendecomposition (4) and consequently also the optimization problem (3), are sensitive to the presence of outlying curves in the sample. Therefore, in order to achieve a robust estimation of the functional PCs, the underlying covariance will be based on a sub sample consisting of the most central data points. The centrality of each observation will here be quantified by an adaptation of a regularized functional Mahalanobis distance for densities.
@@ 3 Notion of Mahalanobis distance for densities
In a traditional multivariate setting, the Mahalanobis distance $M(\textbf{x},\textbf{y};\boldsymbol{\Sigma})$ between two random vectors $\textbf{x},\textbf{y}\in\mathbb{R}^{p}$ with respect to a regular covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{p\times p}$ can be thought of as a Euclidean distance between the two vectors, previously standardized (whitened) by the common covariance
$M^{2}(\textbf{x},\textbf{y};\boldsymbol{\Sigma})$ $=(\textbf{x}-\textbf{y})^{\prime}\boldsymbol{\Sigma}^{-1}(\textbf{x}-\textbf{y})=\|\boldsymbol{\Sigma}^{-1/2}\textbf{x}-\boldsymbol{\Sigma}^{-1/2}\textbf{y}\|^{2}.$
In that manner, the corresponding Mahalanobis norm of a random vector, i.e. the Mahalanobis distance of that random vector from its mean with respect to its covariance, equals the Euclidean norm of that random vector when standardized by its mean and the covariance. That is, $M(\textbf{x};\boldsymbol{\mu},\boldsymbol{\Sigma}):=M(\textbf{x},\boldsymbol{\mu};\boldsymbol{\Sigma})=\|\textbf{x}_{\mathrm{st}}\|$, where $\textbf{x}_{\mathrm{st}}:=\boldsymbol{\Sigma}^{-1/2}(\textbf{x}-\boldsymbol{\mu})$ is a solution to the linear problem $\boldsymbol{\Sigma}^{1/2}\textbf{x}_{\mathrm{st}}=\textbf{x}-\boldsymbol{\mu}$ – for more details on whitening of multivariate data see e.g. (Friedman, 1987; Oja et al., 2016; Kessy et al., 2018).
@@
Oguamalam et al. (2024) show that the $\alpha$-Mahalanobis distance is the Euclidean norm of the solution to (5). Regularization in (5) is also known as $L^{2}$-regularization due to penalizing the $L^{2}$-norm of the solution and is perhaps the most widely used penalization form. However, it is often advantageous to consider regularization using a general, non-identity regularization operator $L:L^{2}(I)\to L^{2}(I)$; see e.g. Morigi et al. (2007); Wang (2012).
@@ 3.1 Regularized standardization of density data
As a corollary of Proposition 3.1, we obtain the explicit solution to (6) for $L=I$.
Corollary 3.1
Let $X_{\mathrm{st}}^{{\alpha},{I}}=_{B}\operatorname*{arg\,min}_{Y\in\mathcal{B}^{2}(I)}\left\{\|{C}_{B^{2},X}^{1/2}Y\ominus X\|_{\mathcal{B}}^{2}+\alpha\|Y\|_{\mathcal{B}}^{2}\right\},$ for regularization parameter $\alpha>0$. Then
$X_{\mathrm{st}}^{\alpha,I}=_{B}\bigoplus_{i=1}^{\infty}\frac{\lambda_{i}^{1/2}}{\lambda_{i}+\alpha}\odot\langle\zeta_{i},X\rangle_{B^{2}}\odot\zeta_{i},$
where $(\lambda_{i},\zeta_{i})$ is an $i$-th eigenpair of $C_{B^{2},X}$ satisfying $C_{B^{2},X}\zeta_{i}=\lambda_{i}\odot\zeta_{i}$, $i\geq 1$.
@@
Choosing regularization operators to be orthogonal projectors, as discussed in the work of Morigi et al. (2007), can be particularly advantageous when we have prior knowledge of certain features or structures within the data. A key motivation for using orthogonal projection regularization operators is that it allows us to directly control the components of the solution within the range and null space of $L$. By projecting onto a subspace that aligns with known properties of the data, we can tailor the regularization to preserve essential characteristics. The following proposition formalizes these insights and provides further motivation for selecting the appropriate regularization strategy by extending the findings of Morigi et al. (2007) (Theorem 2.3) to optimization (6).
@@
where for $k\geq 0$, such that $\lambda_{k}>\lambda_{k+1}$, $X_{1:k}:=\bigoplus_{i=1}^{k}\langle X,\zeta_{i}\rangle_{B^{2}}\otimes\zeta_{i}$ is the best rank-k approximation of $X$, and $X_{(k+1):\infty}:=\bigoplus_{i>k}\langle X,\zeta_{i}\rangle_{B^{2}}\odot\zeta_{i}$ the residual. Additionally, it is often argued that the projection of the data onto the first few principal components is an effective way to retain the most significant features of the original dataset (Ramsay and Silverman, 1997; Hron et al., 2016). Consider, therefore, the projection of the data onto the span of the first $k$ eigenfunctions of $C_{B^{2},X}$ to be the data feature we would like to preserve. As $C_{B^{2},X}^{1/2}X_{1:k}\in\mathrm{span}\{\zeta_{1},\dots,\zeta_{k}\}$, by taking $L$ to be the orthogonal projection onto the complement of $\mathrm{span}\{\zeta_{1},\dots,\zeta_{k}\}$, the regularized approximation $C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{\alpha,L}$ of $X$ is of the form $C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{\alpha,L}=X_{1:k}\oplus C_{B^{2},X}^{1/2}X_{2,\mathrm{st}}^{\alpha,L}$, where $X_{2,\mathrm{st}}^{\alpha,L}\in\mathrm{span}\{\zeta_{k+1},\zeta_{k+2},\dots\}$ is the $L^{2}$-regularized standardization of $X_{(k+1):\infty}$.
Thus, denoting ${W}(k)$ a projection onto the complement of the space spanned by the first $k\geq 0$ eigenfunctions of $C_{B^{2},X}$, hereinafter we consider standardization of the form
@@
Let $X\in\mathcal{B}^{2}(I)$ be a constant mean random density with Bayes-covariance $C_{B^{2},X}$ and let $(\lambda_{i},\zeta_{i})$ be its $i$th eigenpair; $C_{B^{2},X}\zeta_{i}=\lambda_{i}\odot\zeta_{i}$, $i\geq 1$ For $k\geq 0$, such that $\lambda_{k}>\lambda_{k+1}$ let $W(k)$ be the projector onto the complement of the space spanned by the first $k$ eigenfunctions $\zeta_{1},\dots,\zeta_{k}$ of $C_{B^{2},X}$. Then there exists a unique solution $X_{\mathrm{st},C_{B^{2},X}}^{{\alpha},k}$ of (7). Additionally,

### v3

# Robust functional PCA for relative data

@@ Abstract
This paper introduces a robust approach to functional principal component analysis (FPCA) for relative data, particularly density functions. While recent papers have studied density data within the Bayes space framework, there has been limited focus on developing robust methods to effectively handle anomalous observations and large noise. To address this, we extend the Mahalanobis distance concept to Bayes spaces, proposing its regularized version that accounts for the constraints inherent in density data. Based on this extension, we introduce a new method, robust density principal component analysis (RDPCA), for more accurate estimation of functional principal components in the presence of outliers. The method’s performance is validated through simulations and real-world applications, showing its ability to improve covariance estimation and principal component analysis compared to traditional methods.
@@ 1 Introduction
Functional data analysis (FDA) has developed rapidly over the past decade (e.g., Ramsay and Silverman (2005)), yet many applied settings involve relative functional data—functions where the key information lies not in their absolute magnitude, but the dynamics or distributional pattern of the measured quantity carries relevant information. Examples include age-specific fertility or mortality rates, spectral curves in materials science, and normalized sensor signals, where data are typically rescaled and interpreted in terms of relative variation across a domain.
These data often share structural features with probability density functions (PDFs)—non-negativity, unit integral, and a relative (compositional) structure. While PDFs are a motivating case, the broader goal is to analyze functions where the shape or internal distribution carries the signal of interest, and scale is irrelevant or removed. For the simplicity of the notation, we refer to such functional data as densities, emphasizing that the probability density functions are just one example of the functional data we consider in this context.
To analyze such data, we use compositional data analysis (CoDA) extended to infinite dimensions, modeling functions as infinite-dimensional compositions in a Bayes space, a Hilbert space respecting their relative nature (Egozcue et al., 2006; van den Boogaart et al., 2010). Within this framework, simplicial functional PCA (SFPCA) (Hron et al., 2016) offers a dimension reduction tool tailored to relative functional data. However, SFPCA depends on the sample covariance operator and is sensitive to outliers, limiting its robustness in practice. To address this, robust alternatives using trimmed covariance estimators and center-outward orderings have been proposed. Depth measures such as spatial depth (Menafoglio et al., 2021) and others (Fraiman and Muniz, 2001; Febrero et al., 2008; López-Pintado and Romo, 2009; Arribas-Gil and Romo, 2014; Chakraborty and Chaudhuri, 2014) offer one route, while Mahalanobis-type distances—widely used in multivariate and functional settings (Secchi et al., 2013; Galeano et al., 2015; Ghiglietti et al., 2017; Berrendero et al., 2020)—provide another, such as the minimum regularized covariance trace (MRCT) estimator (Oguamalam et al., 2024). In this paper, we focus on the latter group of methods, extending the concept of regularized Mahalanobis distance for univariate functional data (later called RDMD) to the Bayes space. We then use this notion to construct a robust covariance estimator and a corresponding robust PCA method tailored for densities.
@@
The rest of the paper is organized as follows: Section 2 discusses the concept of Bayes spaces and SFPCA and showcases potential problems that the outliers in the sample can cause. In Section 3 we propose a way of standardizing (whitening) relative functional data, thus extending and generalizing the notions of truncated- and $\alpha$-Mahalanobis distance from Galeano et al. (2015) and Berrendero et al. (2020), respectively, to Bayes Hilbert spaces. In Section 4, the novel, robust density PCA (RDPCA) method is introduced. Its performance is evaluated in a simulation study and real-data examples in Sections 5 and 6, respectively. Proofs of technical results are given in Supplement Section LABEL:supp:proofs.
@@ 2.1 Bayes space
If we expand the horizon to infinite-dimensional compositions, we consider positive functions that are typically represented with unit integral constraint. Intuitively, such compositions can be thought of as density functions. Therefore, statistical concepts designed for unconstrained functional data need to be adapted, as they do not account for this feature, leading to potentially meaningless results when applied to densities (Delicado, 2011; Hron et al., 2016; Menafoglio et al., 2018). Consequently, to work with functional compositions, it is necessary to establish the fundamental concepts of the underlying space structure.
The foundational results were initially presented by Egozcue et al. (2006) and were subsequently extended by van den Boogaart et al. (2010), van den Boogaart et al. (2014), and Talská et al. (2020). The framework for probability density functions is based on a measurable space $(\Omega,\mathcal{A})$, typically set to $(\mathbb{R},\mathfrak{B}(\mathbb{R}))$, with $\mathfrak{B}(\mathbb{R})$ as the Borel $\sigma$-algebra on $\mathbb{R}$. Within this space, we consider the set of measures $\mu$ that are absolutely continuous with respect to a reference measure $\lambda$, here assumed to be the Lebesgue measure with compact support $I=[a,b]\subset\mathbb{R}$ and with length $\eta:=b-a$. While the authors of van den Boogaart et al. (2014) developed the theory of Bayes space for more general measures $\lambda$, in practice, commonly the Lebesgue measure is chosen. Furthermore, usually the measures $\mu$ are identified with their respective Radon-Nikodym derivative, density $f$. For these densities, a notion of equivalence in the Bayes sense for multiples is introduced, i.e $f=_{B}g$ if $f=c\cdot g$ for some $c\in\mathbb{R}.$ W.l.o.g. within these $(=_{B})$-equivalence classes, we take the element satisfying the unit integral constraint as the representative.
To equip the sample space of these density functions with a structure of a Hilbert space on a given bounded domain, Egozcue et al. (2006) followed the ideas grounding the Aitchison geometry, and extended its operations $(\oplus,\odot)$ (a.k.a. perturbation and powering), and the inner product $\langle\cdot,\cdot\rangle$ to the infinite-dimensional setting by defining
$(f\oplus g)(t)=_{B}f(t)g(t),\quad(\alpha\odot f)(t)=_{B}(f(t))^{\alpha},\quad\langle f,g\rangle_{B^{2}}=\frac{1}{2\eta}\int_{I}\int_{I}\text{log}\frac{f(t)}{f(s)}\text{log}\frac{g(t)}{g(s)}\mathrm{d}t\,\mathrm{d}s,$
for densities $f,g$ and $\alpha\in\mathbb{R}$. In this context, negative perturbation $(\ominus)$ is defined as $f\ominus g=f\oplus(-1\odot g)=_{B}f/g$. Endowing the subspace $\mathcal{B}^{2}(I)$ of positive densities on $I$ with a square-integrable logarithm (van den Boogaart et al., 2014) with this structure yields a separable Hilbert space. In practice, the centered logratio (clr) transformation
$\text{clr}(f)=\text{log}(f)-\frac{1}{\eta}\int_{I}(\text{log}(f))(t)\mathrm{d}t$ (1)
@@
$\langle f,g\rangle_{B^{2}}=~\langle\text{clr}(f),\text{clr}(g)\rangle_{L^{2}}~~\text{and}~~\text{clr}((\alpha\odot f)\oplus(\beta\odot g))=~\alpha\cdot\text{clr}(f)+\beta\cdot\text{clr}(g),$
making clr transformation an isometry.
@@ 2.2 Moments of random densities
for every $f,g\in\mathcal{B}^{2}(I)$. We write $C_{B^{2},X}=\mathbb{E}((X\ominus\mu_{B^{2},X})\otimes_{B^{2}}(X\ominus\mu_{B^{2},X})),$ where for $f,g\in\mathcal{B}^{2}(I)$ the outer product $f\otimes_{B^{2}}g:\mathcal{B}^{2}(I)\to\mathcal{B}^{2}(I)$ is the linear operator satisfying $(f\otimes_{B^{2}}g)(h)=\langle g,h\rangle_{B^{2}}f$, for $h\in\mathcal{B}^{2}(I)$. We require the density variable $X$ to possess finite fourth moment, i.e., $\mathbb{E}[\|X\|^{4}_{B^{2}}]<\infty$. Under this assumption, the covariance operator is a symmetric positive-definite Hilbert-Schmidt operator and admits the eigen-decomposition
$C_{B^{2},X}f=\sum_{j=1}^{+\infty}\lambda_{j}\langle f,\xi_{j}\rangle\xi_{j},$
where $(\lambda_{j},\xi_{j})$ are the eigenpairs of $C_{B^{2},X}$, conventionally ordered as $\lambda_{1}\geq\lambda_{2}\geq...\geq 0$. Note that all the eigenvalues are non-negative and, under the conditions above, $C_{B^{2},X}$ is trace-class, i.e., $\sum_{j=1}^{+\infty}\lambda_{j}<\infty$ (see, e.g., Horváth and Kokoszka (2012)). Lemma 2.1 gives the connection between the Bayes mean density $\mu_{B^{2},X}$ and covariance $C_{B^{2},X}$ and the “traditional" $L^{2}$-mean and the covariance of $\mathrm{clr}(X)$. The proof of Lemma 2.1 can be found in Section LABEL:supp:proofs of the supplement.
@@ 2.3 Simplicial functional PCA
As in the multivariate principal component analysis, functional PCA aims to search for the main modes of variability within a dataset, see e.g. Ramsay and Silverman (2005). More recently, this concept was extended to the Bayes spaces setting in Hron et al. (2016). The authors coined the term simplicial functional PCA, referring to the simplex, which is the multivariate counterpart of the Bayes space. More precisely, given a data set $X_{1},\ldots,X_{n}\in\mathcal{B}^{2}(I)$, the problem of finding the first principal function $\zeta_{1}$ is defined as
[…]
@@
However, as previously emphasized, the eigendecomposition (4) and consequently also the optimization problem (3), are sensitive to the presence of outlying curves in the sample. Therefore, in order to achieve a robust estimation of the functional PCs, the underlying covariance will be based on a sub-sample consisting of the most central data points. The centrality of each observation will here be quantified by an adaptation of a regularized functional Mahalanobis distance for densities. For more insight into the notion of centrality/outlyingness for densities, see Remark 2.1.
Remark 2.1
In contrast to traditional FDA settings, identifying atypical curves in Bayes spaces requires rethinking outlier concepts. Because relative functions are compositional, local deviations affect the entire curve, and scaling, a key factor in conventional magnitude outliers, has no effect. Thus, outliers in this context reflect structural (shape) anomalies, such as unexpected redistribution of mass or signal across the domain. We detect such curves using the proposed concept of robust, regularized MD, flagging outliers based on the chosen quantile of its limiting distribution. We emphasize, however, that the goal of the present work is not to develop a method for outlier detection or to provide a taxonomy of outliers in relative functional data. Rather, we propose a novel approach to constructing robust PCA methods for relative functional data.
@@ 3 Notion of Mahalanobis distance for densities
In a traditional multivariate setting, the Mahalanobis distance $M(\textbf{x},\textbf{y};\boldsymbol{\Sigma})$ between two random vectors $\textbf{x},\textbf{y}\in\mathbb{R}^{p}$ with respect to a regular covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{p\times p}$ can be thought of as a Euclidean distance between the two vectors, previously standardized (whitened) by the common covariance $M^{2}(\textbf{x},\textbf{y};\boldsymbol{\Sigma})=(\textbf{x}-\textbf{y})^{\prime}\boldsymbol{\Sigma}^{-1}(\textbf{x}-\textbf{y})=\|\boldsymbol{\Sigma}^{-1/2}\textbf{x}-\boldsymbol{\Sigma}^{-1/2}\textbf{y}\|^{2}.$ In that manner, the corresponding Mahalanobis norm of a random vector, i.e. the Mahalanobis distance of that random vector from its mean with respect to its covariance, equals the Euclidean norm of that random vector when standardized by its mean and the covariance. That is, $M(\textbf{x};\boldsymbol{\mu},\boldsymbol{\Sigma}):=M(\textbf{x},\boldsymbol{\mu};\boldsymbol{\Sigma})=\|\textbf{x}_{\mathrm{st}}\|$, where $\textbf{x}_{\mathrm{st}}:=\boldsymbol{\Sigma}^{-1/2}(\textbf{x}-\boldsymbol{\mu})$ is a solution to the linear problem $\boldsymbol{\Sigma}^{1/2}\textbf{x}_{\mathrm{st}}=\textbf{x}-\boldsymbol{\mu}$; for more details on whitening of multivariate data see e.g. (Friedman, 1987; Oja et al., 2016; Kessy et al., 2018).
@@
Oguamalam et al. (2024) show that the $\alpha$-Mahalanobis distance is the Euclidean norm of the solution to (5). Regularization in (5) is also known as $L^{2}$-regularization due to penalizing the $L^{2}$-norm of the solution and is perhaps the most widely used penalization form. However, it is often advantageous to consider regularization using a more general regularization operator $L:L^{2}(I)\to L^{2}(I)$; see e.g. Morigi et al. (2007); Wang (2012).
@@ 3.1 Regularized standardization of density data
Corollary LABEL:cor:L_2_standardization_explicit_form in the Supplement LABEL:supp:aux_results gives the explicit solution to (6) for $L=I$.
@@
Choosing regularization operators to be orthogonal projectors, as discussed in the work of Morigi et al. (2007), can be particularly advantageous when we have prior knowledge of certain features or structures within the data. A key motivation for using orthogonal projection regularization operators is that they allow us to directly control the components of the solution within the range and null space of $L$. By projecting onto a subspace that aligns with known properties of the data, we can tailor the regularization to preserve essential characteristics. The following proposition formalizes these insights and provides further motivation for selecting the appropriate regularization strategy by extending the findings of Morigi et al. (2007) (Theorem 2.3) to optimization (6).
@@
where for $k\geq 0$, such that $\lambda_{k}>\lambda_{k+1}$, $X_{1:k}:=\bigoplus_{i=1}^{k}\langle X,\zeta_{i}\rangle_{B^{2}}\odot\zeta_{i}$ is the best rank-k approximation of $X$, and $X_{(k+1):\infty}:=\bigoplus_{i>k}\langle X,\zeta_{i}\rangle_{B^{2}}\odot\zeta_{i}$ the residual. Additionally, it is often argued that the projection of the data onto the first few principal components is an effective way to retain the most significant features of the original dataset (Ramsay and Silverman, 1997; Hron et al., 2016). Consider, therefore, the projection of the data onto the span of the first $k$ eigenfunctions of $C_{B^{2},X}$ to be the data feature we would like to preserve. As $C_{B^{2},X}^{1/2}X_{1:k}\in\mathrm{span}\{\zeta_{1},\dots,\zeta_{k}\}$, by taking $L$ to be the orthogonal projection onto the complement of $\mathrm{span}\{\zeta_{1},\dots,\zeta_{k}\}$, the regularized approximation $C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{\alpha,L}$ of $X$ is of the form $C_{B^{2},X}^{1/2}X_{\mathrm{st}}^{\alpha,L}=X_{1:k}\oplus C_{B^{2},X}^{1/2}X_{2,\mathrm{st}}^{\alpha,L}$, where $X_{2,\mathrm{st}}^{\alpha,L}\in\mathrm{span}\{\zeta_{k+1},\zeta_{k+2},\dots\}$ is the $L^{2}$-regularized standardization of $X_{(k+1):\infty}$.
Thus, denoting W(k){W}(k) a projection onto the complement of the space spanned by the first $k\geq 0$ eigenfunctions of $C_{B^{2},X}$, hereinafter we consider standardization of the form
@@
Let $X\in\mathcal{B}^{2}(I)$ be a constant mean random density with Bayes-covariance $C_{B^{2},X}$ and let $(\lambda_{i},\zeta_{i})$ be its $i$th eigenpair; $C_{B^{2},X}\zeta_{i}=\lambda_{i}\odot\zeta_{i}$, $i\geq 1$ . For $k\geq 0$, such that $\lambda_{k}>\lambda_{k+1}$ let $W(k)$ be the projector onto the complement of the space spanned by the first $k$ eigenfunctions $\zeta_{1},\dots,\zeta_{k}$ of $C_{B^{2},X}$. Then there exists a unique solution $X_{\mathrm{st},C_{B^{2},X}}^{{\alpha},k}$ of (7). Additionally,